# GelSLAM: A Real-time, High-Fidelity, and Robust 3D Tactile SLAM System

### v1

@@ Abstract
Accurately perceiving an object’s pose and shape is essential for precise grasping and manipulation. Compared to common vision-based methods, tactile sensing offers advantages in precision and immunity to occlusion when tracking and reconstructing objects in contact. This makes it particularly valuable for in-hand and other high-precision manipulation tasks. In this work, we present GelSLAM, a real-time 3D SLAM system that relies solely on tactile sensing to estimate object pose over long periods and reconstruct object shapes with high fidelity. Unlike traditional point cloud–based approaches, GelSLAM uses tactile-derived surface normals and curvatures for robust tracking and loop closure. It can track object motion in real time with low error and minimal drift, and reconstruct shapes with submillimeter accuracy, even for low-texture objects such as wooden tools. GelSLAM extends tactile sensing beyond local contact to enable global, long-horizon spatial perception, and we believe it will serve as a foundation for many precise manipulation tasks involving interaction with objects in hand. The video demo is available on our website: https://joehjhuang.github.io/gelslam.
@@ I Introduction
In this work, we aim to achieve real-time long-horizon object tracking and high-fidelity object 3D reconstruction using tactile input alone. Object tracking in a hand is particularly useful for robotic manipulation, where vision is often occluded during contact, while object 3D reconstruction benefits applications in AR/VR, biology, archaeology, dentistry, and geology that require detailed geometric analysis. We focus on GelSight-related tactile sensors [23, 60], which are known for their high spatial resolution for capturing local surface geometry.
@@
To the best of our knowledge, GelSLAM is the first tactile-only system to achieve object-level 3D reconstruction with this degree of robustness, accuracy, and fidelity. GelSLAM shows that what was once a local sensing modality, touch, can move beyond its inherent limitations to support global, long-horizon spatial understanding. We believe it lays a foundation for future research in tactile-based perception, visuo-tactile integration, contact-rich manipulation, and dexterous manipulation. Beyond robotics, potential applications include dental scanning, biological phenotype analysis, in-the-wild geological surface inspection, archaeological reconstruction, and detailed shape capture for AR/VR. To support continued progress in the field, we will release the full source code and dataset to the community.
@@ II-A Object Pose Estimation
Despite its popularity in pose estimation, vision often struggles in real-world robotics due to occlusions [49] and poor performance on optically challenging objects [7]. Tactile sensing, on the other hand, offers high precision and is immune to occlusions as well as transparent or specular surfaces. When the object’s geometry is known, [4] estimates a grasped object’s 6DoF pose by using a PCA-based covariance matching method to align the statistical structure of tactile array data with the local object geometry. Using a vision-based tactile sensor, DIGIT [22], Midas-Touch [47] applies a particle filter to track the evolving 6DoF pose distribution of a fingertip sliding across a known object’s surface. Tac2Pose [3] estimates the pose of a known object from GelSlim [50] tactile readings by matching the sensed contact shape to simulated shapes in an object-specific, contrastively learned embedding. In robot manipulation, a more common scenario is handling previously unseen objects. Wang et al. [54] proposed estimating object transformations by applying the Iterative Closest Point (ICP) algorithm to tactile-derived point clouds. Similarly, PatchGraph [45] estimates frame-to-frame transformations with ICP and incorporates motion and geometric constraints in a factor graph for more precise object pose tracking. Finally, NormalFlow [12] uses tactile-derived normal maps to achieve very accurate pose tracking and performs well even on low-textured objects.
@@ III-C1 Surface Normal Map
For auxiliary purposes, the height map $\mathbf{H}_{i}\in\mathbb{R}^{H\times W}$ is computed by integrating the predicted surface gradients using a 2D fast Poisson solver [60]. The contact mask $\mathbf{C}_{i}\in\{0,1\}^{H\times W}$ is then derived by thresholding the height map and the RGB value changes of each pixel in the tactile image. The height map can be converted into the point cloud representation by placing each pixel at its image location with the height value as its z-coordinate. The tactile-derived mesh can then be obtained by triangulating neighboring points in the point cloud according to the pixel grid structure.
@@ IV-B3 Pose Graph Optimization
IV-B4 GNC Pose Graph Optimization (Optional)
In some cases, objects with repetitive surface patterns may cause loop detection to incorrectly associate frames from different locations. Such repetition is uncommon in GelSight tactile sensing due to its high spatial resolution, but it can occur in industrially manufactured objects. Our least-squares optimization in Equation (3) is sensitive to these false matches, as they introduce outliers. To address this, we apply the GNC solver [58] to robustly optimize the pose graph by rejecting outlier constraints. However, for objects without repetitive patterns, this method may misclassify true matches as outliers. It also incurs much higher computational cost, making it suitable only for offline reconstruction. Therefore, we default to the regular pose graph optimization (Section IV-B3) and only apply the GNC solver when failures arise due to repetitive patterns.
@@ V Experiment: Long-horizon Tracking
In this section, we evaluate the performance of GelSLAM on long-horizon 6DoF object pose tracking. Our experiments use the GelSight Mini tactile sensor, configured to a resized resolution of $320\times 240$. The sensor has an effective sensing area of $20\,\text{mm}\times 15\,\text{mm}$, a frame rate of 25 Hz, and a maximum indentation depth of approximately $2$ mm. The system is evaluated on an AMD Ryzen 7 PRO 7840U CPU with 8 cores at 3.3 GHz. GelSLAM is implemented in ROS2 [33].
@@ V-B System Characteristics and Performance
Frames Keyframes
Coverage
Keyframes
Loop
Candidates
Loops
@@
Table III presents the runtime breakdown of each GelSLAM module during online operation. The tracking module runs faster than the GelSight’s 25 Hz frame rate. The method “compute geometric properties” refers to the step that takes a GelSight image $\mathbf{F}_{t}$ and computes $\{\mathbf{I}_{t},\mathbf{L}_{t},\mathbf{H}_{t},\mathbf{C}_{t}\}$. In the loop closure module, loop detection is the most time-consuming step. Its runtime varies significantly across episodes, as it depends heavily on the number of loop candidates. The module is triggered only when a new keyframe is inserted, which occurs every several frames, making the runtime more than sufficient. Occasionally, if the module takes too long, a keyframe may be skipped for loop closure. This does not affect system stability but may slightly reduce accuracy by missing some loop constraints.
@@
Loop
Closure
Loop Detection 12.6 25.2
@@ V-D Long-horizon Tracking Results
We organize our ablation study in Table IV around three key components:
Keyframe selection method: Without loop closure, using our keyframe selection method (NF-KF) achieves better results than the baseline without keyframes (NF). Introducing keyframes helps reduce drift by avoiding dense frame-to-frame composition and also lowers the computation in the loop closure module. Compared to the earlier keyframe selection method (NF-OrigKF) [12], our keyframe selection method achieves similar performance while running twice as fast, requiring only a single NormalFlow computation per frame instead of two.
Loop detection method: Incorporating loop closure using SIFT matching alone (GS-OnlySIFT) does not improve performance relative to our keyframe-based tracker (NF-KF) and can even degrade it due to outlier loops. Using NormalFlow alone for loop detection (GS-OnlyNF) achieves better results because our NormalFlow failure detection mechanism can reject incorrect loop closures. However, it is significantly slower, as it must compare all pairs without SIFT-based pre-selection, which is about 40 times faster than NormalFlow. Moreover, NormalFlow without a good initialization can easily get stuck in local minima and falsely reject valid loops. The best results come from the two-stage pipeline (GS-Offline), where SIFT matching provides a reliable initialization, enabling NormalFlow to detect loops it would otherwise miss.
SIFT feature source: We also study how the source of SIFT features affects robustness. Applying SIFT to the raw GelSight image (GS-ImageSIFT) is less reliable due to the image’s sensitivity to spatial transformations. In contrast, using the curvature map (GS-Offline) yields better results by improving SIFT matching to provide more reliable NormalFlow initializations. Note that, unlike GS-OnlySIFT, GS-ImageSIFT includes the NormalFlow refinement stage for loop detection.
Figure 8 compares tracking error over time for NormalFlow (NF), NormalFlow with Keyframes (NF-KF), and our method across two episodes. Our approach accurately identifies many loops even under limited texture and effectively corrects drift, leading to lower tracking error than the baselines.
@@ VI-A Data Collection
We collect one GelSight video per object for qualitative analysis on a set of 15 real-world objects, including $3$ tool handles, $6$ common food items, $4$ rocks and fossils, and $2$ textured objects. The smallest object is Seed ($8\,mm\times 8\,mm\times 8\,mm$), and the largest is Avocado ($85\,mm\times 61\,mm\times 58\,mm$). All objects have at least some surface texture, with the smoothest being the handle of a wooden spoon. Video durations range from $1$ to $30$ minutes. Fig. 12 shows each object with its size and video duration. For quantitative analysis, we use 5 object models with known ground-truth geometries. Each model is 3D-printed in two sizes using a Form 3+ printer, with a reported average dimensional deviation of up to 0.1 mm, resulting in a total of 10 printed objects. The smallest and largest objects are a Seed and Lime with a diameter of about 9 mm and 30 mm, respectively. All data are collected in an “in-the-wild” setup, where we manually hold both the object and the GelSight Mini sensor (Fig. 9) during scanning. Contact breaks and re-initializations are common during data collection, sometimes exceeding 100 occurrences per object. Scanning paths are guided by real-time reconstruction feedback, which may become laggy after approximately $10$ minutes due to the growing size of the displayed mesh. In such cases, we perform multiple shorter scans and merge the resulting videos into a single sequence for reconstruction.
@@ VI-B System Characteristics and Performance
We run GelSLAM in offline mode on all object videos. Since the objects provide sufficient texture, we apply fixed and less conservative thresholds for NormalFlow failure detection: the estimate is rejected if $\text{CCS}<0.7$ or $\text{SCR}<0.3$ across all objects. Table V summarizes reconstruction statistics for four objects with varying video lengths, and Fig. 10 shows their corresponding pose graphs.
@@
Object Name Avocado
Large
Rock
Peanut Seed
@@
GelSLAM Offline
Runtime (min)
40.92 29.82 4.78 0.6
@@ VI-C Qualitative Results
Fig. 12 shows tactile-based object-level 3D reconstructions for the $15$ real-world objects. All loops detected by GelSLAM are correct except for those in the Carved Bead, which contains repetitive patterns. These incorrect loops, caused by the repetitive patterns, can not be filtered using local information alone. Therefore, we apply GNC pose graph optimization (Section IV-B4) to the Carved Bead, while all other objects use the regular optimization method (Section IV-B3). The results demonstrate that GelSLAM correctly relates GelSight images spatially and globally reconstructs objects across a wide range of textures and sizes, even on low-texture objects such as the wooden spoon handle. The reconstructed geometry also exhibits remarkable detail. Fig. 14 compares regular and GNC pose graph optimization on objects with and without repetitive textures. For objects with distinct textures, GNC optimization may incorrectly reject valid loops, resulting in a locally disconnected mesh. In contrast, for objects with repetitive textures, outlier loops are unavoidable, and even a few can cause catastrophic failure in the regular optimization method. GNC optimization, however, effectively filters these outliers and significantly improves reconstruction quality.
@@
Figure 14: Comparison of regular and GNC pose graph optimization. On objects without repetitive textures, GNC may mistakenly reject valid loops, leading to locally disconnected reconstruction. For objects with repetitive textures, outlier loops are unavoidable, and GNC improves reconstruction by filtering them.

### v2

@@ Abstract
Accurately perceiving an object’s pose and shape is essential for precise grasping and manipulation. Compared to common vision-based methods, tactile sensing offers advantages in precision and immunity to occlusion when tracking and reconstructing objects in contact. This makes it particularly valuable for in-hand and other high-precision manipulation tasks. In this work, we present GelSLAM, a real-time 3D SLAM system that relies solely on tactile sensing to estimate object pose over long periods and reconstruct object shapes with high fidelity. Unlike traditional point cloud–based approaches, GelSLAM uses tactile-derived surface normals and curvatures for robust tracking and loop closure. It can track object motion in real time with low error and minimal drift, and reconstruct shapes with submillimeter accuracy, even for low-texture objects such as wooden tools. GelSLAM extends tactile sensing beyond local contact to enable global, long-horizon spatial perception, and we believe it will serve as a foundation for many precise manipulation tasks involving interaction with objects in hand. The video demo, code, and dataset are available at https://joehjhuang.github.io/gelslam.
@@ I Introduction
In this work, we aim to achieve real-time long-horizon object tracking and high-fidelity object 3D reconstruction using tactile input alone. Real-time object tracking using touch is particularly useful for contact-rich manipulation with robot hands, where high precision is required but vision is often occluded or suffers from low lighting. Meanwhile, object 3D reconstruction using tactile sensing benefits applications in biology, archaeology, dentistry, and geology that require capturing fine surface details often missed by vision. We focus on GelSight-related tactile sensors [23, 62], which are known for their high spatial resolution for capturing local surface geometry.
@@
Contributions: Our main contribution is the three modules, each incorporating novel components tailored to tactile sensing, as well as the well-engineered system that integrates them into a robust tactile SLAM framework. The novel tactile-oriented components include tracking failure detection, keyframe selection, and robust loop detection. Together, these components enable GelSLAM to operate reliably in real time over tens of thousands of frames, whereas prior methods scale only to hundreds of frames and fail frequently. To the best of our knowledge, GelSLAM is the first tactile-only system to achieve object-level 3D reconstruction with this degree of robustness, accuracy, and fidelity. GelSLAM shows that what was once a local sensing modality, touch, can move beyond its inherent limitations to support global, long-horizon spatial understanding. We believe it lays a foundation for future research in tactile-based perception, visuo-tactile integration, contact-rich manipulation, and dexterous manipulation. Beyond robotics, potential applications include dental scanning, biological phenotype analysis, in-the-wild geological surface inspection, archaeological reconstruction, and detailed shape capture for AR/VR. To support continued progress in the field, we release the full source code and dataset to the community.
@@ II-A Object Pose Estimation
Despite its popularity in pose estimation, vision often struggles in real-world robotics due to occlusions [51] and poor performance on optically challenging objects [7]. Tactile sensing, on the other hand, offers high precision and is immune to occlusions as well as transparent or specular surfaces. When the object’s geometry is known, [4] estimates a grasped object’s 6DoF pose by using a PCA-based covariance matching method to align the statistical structure of tactile array data with the local object geometry. Using a vision-based tactile sensor, DIGIT [22], Midas-Touch [49] applies a particle filter to track the evolving 6DoF pose distribution of a fingertip sliding across a known object’s surface. Tac2Pose [3] estimates the pose of a known object from GelSlim [52] tactile readings by matching the sensed contact shape to simulated shapes in an object-specific, contrastively learned embedding. In robot manipulation, a more common scenario is handling previously unseen objects. Wang et al. [56] proposed estimating object transformations by applying the Iterative Closest Point (ICP) algorithm to tactile-derived point clouds. NormalFlow [12] uses tactile-derived normal maps to achieve very accurate pose tracking and performs well even on low-textured objects. Another common practice is to incorporate motion and geometric priors within a factor-graph framework, together with frame-to-frame relational constraints, to improve pose tracking accuracy. Sodhi et al. [45] obtain these constraints from learned relative pose measurements between tactile image pairs. PatchGraph [46] follows a similar formulation but estimates frame-to-frame transformations using ICP. Suresh et al. [48] instead derive them from implicit-surface contact constraint, where the implicit surface is obtained by integrating contact information over time.
@@ III-C1 Surface Normal Map
For auxiliary purposes, the height map $\mathbf{H}_{i}\in\mathbb{R}^{H\times W}$ is computed by integrating the predicted surface gradients using a 2D fast Poisson solver [62]. Specifically, the solver formulates gradient integration as a least-squares optimization problem that solves for the height map whose spatial derivatives best match the predicted gradients. The contact mask $\mathbf{C}_{i}\in\{0,1\}^{H\times W}$ is then derived by thresholding the height map and the RGB value changes of each pixel in the tactile image. The height map can be converted into the point cloud representation by placing each pixel at its image location with the height value as its z-coordinate. The tactile-derived mesh can then be obtained by triangulating neighboring points in the point cloud according to the pixel grid structure.
@@ IV-C The Reconstruction Module
IV-D Summary of Contributions
In summary, our main contributions are the algorithmic design of novel SLAM components tailored to tactile sensing and their integration into three modules that form a well-engineered tactile SLAM system. The proposed components include NormalFlow failure detection, keyframe selection, and loop detection. Conventional SLAM components do not transfer well to tactile sensing due to the fundamentally different challenges it poses compared to vision and other sensing modalities. Accordingly, we incorporate our key insight on differential representations into the design of these components. For example, standard loop detection methods based on matching SIFT features assume rotational and translational invariance, which tactile images do not exhibit. Instead, we match SIFT features of the curvature map, which can be accurately estimated from tactile data, are invariant to rigid transformations, and remain informative even when tactile point clouds are nearly planar.
While GelSLAM is built on NormalFlow [12], NormalFlow itself can not detect its own tracking failures or perform relocalization after contact is lost, and remains a purely local method. Our novel components are key to enabling robust, real-time, and large-scale tracking and reconstruction. With these components, GelSLAM robustly detects thousands of loops across tens of thousands of frames, whereas NormalFlow and other existing approaches, such as Tac2Structure [32], scale only to hundreds of frames and suffer from frequent false loop detections.
@@ V Experiment: Long-horizon Tracking
In this section, we evaluate the performance of GelSLAM on real-time, long-horizon 6DoF object pose tracking. Our experiments use the GelSight Mini tactile sensor, configured to a resized resolution of $320\times 240$. The sensor has an effective sensing area of $20\,\text{mm}\times 15\,\text{mm}$, a frame rate of 25 Hz, and a maximum indentation depth of about $2$ mm. The system is evaluated on an AMD Ryzen 7 PRO 7840U CPU with 8 cores at 3.3 GHz. GelSLAM is implemented in ROS2 [33].
@@ V-B System Characteristics and Performance
Frames Keyframes Coverage Keyframes Loop Candidates Loops
@@
Table III presents the runtime breakdown of each GelSLAM module during online operation. The tracking module runs faster than the GelSight’s 25 Hz frame rate. The method “compute geometric properties” refers to the step that takes a GelSight image $\mathbf{F}_{t}$ and computes $\{\mathbf{I}_{t},\mathbf{L}_{t},\mathbf{H}_{t},\mathbf{C}_{t}\}$. In the loop closure module, loop detection is the most time-consuming step, with a runtime ($12.6\pm 25.2$ ms) that depends on the number of loop candidates. Since this module is triggered only when a new keyframe is inserted, on average every $310$ ms, its runtime is well within the time budget for online operation.
As the episode length increases, loop detection runtime can increase, potentially causing some keyframes to be skipped for loop closure during online execution. To evaluate this effect in long episodes, we analyze a seven-minute episode tracking the Hammer. Fig. 7 shows the number of detected and skipped loops over time. No loops are skipped during the first minute, which already exceeds the duration of most manipulation tasks. Over the full seven-minute episode, fewer than 5% of loops are skipped. Overall, the impact on online operation, due to skipped loop closures, is negligible for realistic task durations and remains small even in unusually long episodes.
Figure 7: Detected and skipped loop closures over time during online execution. Fewer than 5% of loops are skipped, even in a very long episode.
@@
Loop Closure Loop Detection 12.6 25.2
@@ V-D Long-horizon Tracking Results
We organize our ablation study in Table IV around three key engineering decisions, also highlighting the impact of differential representations in our SLAM components.
Keyframe selection method: Without loop closure, using our keyframe selection method (NF-KF) achieves better results than the baseline without keyframes (NF). Introducing keyframes helps reduce drift by avoiding dense frame-to-frame composition and also lowers the computation in the loop closure module. Compared to the earlier keyframe selection method (NF-OrigKF) [12], our approach (NF-KF) uses differential-representation-based metrics (CCS and SCR) to achieve similar performance while running twice as fast, requiring only a single NormalFlow computation per frame instead of two.
Loop detection method: Incorporating loop closure using SIFT matching alone (GS-OnlySIFT) does not improve performance relative to our keyframe-based tracker (NF-KF) and can even degrade it due to outlier loops. Using NormalFlow alone for loop detection (GS-OnlyNF) achieves better results because our NormalFlow failure detection mechanism, based on the proposed CCS and SCR metrics, can reject incorrect loop closures. However, it is significantly slower, as it must compare all pairs without SIFT-based pre-selection, which is about 40 times faster than NormalFlow. Moreover, NormalFlow without a good initialization can easily get stuck in local minima and falsely reject valid loops. The best results come from the two-stage pipeline (GS-Offline), where SIFT matching provides a reliable initialization, enabling NormalFlow to detect loops it would otherwise miss.
SIFT feature source: We also study how the source of SIFT features affects robustness. Loop detection using SIFT extracted from the raw GelSight image (GS-ImageSIFT) is less reliable due to the image’s sensitivity to spatial transformations, while SIFT from a differential representation (GS-Offline, using the curvature map) improves matching reliability, reducing rotational tracking error by $34\%$. Note that, unlike GS-OnlySIFT, GS-ImageSIFT includes the NormalFlow refinement stage for loop detection.
V-E Case Studies on Long-Horizon Tracking
@@
Figure 10: Tracking error of our approach averaged over time across CCS and SCR thresholds. Tracking error is high at extreme thresholds but remains low across a wide range, with values much lower than the baselines indicated on the colorbar.
We study two representative episodes (one on Pliers and one on Spatula) to analyze how the tracking error of our approach evolves over time, evaluate its sensitivity to the CCS and SCR thresholds, and examine its relocalization capability. Fig. 9 compares tracking error over time for NormalFlow (NF), NormalFlow with keyframes (NF-KF), and our method. Despite limited surface texture, our approach reliably identifies many loops and corrects accumulated drift, achieving consistently lower tracking error than the baselines.
Fig. 10 illustrates the sensitivity of our approach to the CCS threshold $(\in[-1,1])$ and SCR threshold $(\in[0,1])$. Setting either threshold to 1 makes every frame a keyframe and rejects all loop closures, causing GelSLAM to degenerate to NormalFlow. Across a wide range of CCS and SCR thresholds, the tracking error remains low, showing that performance is insensitive to the exact threshold choice. When the CCS threshold is too low (below 0), accuracy degrades due to accepting false loop closures; in practice, objects with less surface texture require a higher CCS threshold to avoid this issue. For long-horizon tracking, we find that a CCS threshold of 0.85 and an SCR threshold of 0.3 perform well across most everyday objects.
We evaluate the relocalization capability of our method by simulating periodic contact loss. Specifically, every 4 seconds (100 frames), we remove a 1-second segment (25 frames) from the sequence. Fig. 11 compares the tracking results of our approach obtained using the full episodes with those obtained under simulated contact removals. Once contact is re-established, our method reliably relocalizes and achieves tracking performance nearly identical to that obtained without contact removal, although relocalization may be delayed when contact resumes far from previously contacted regions.
Figure 11: Tracking error comparison of our approach with and without simulated contact loss. Our method can reliably relocalize after contact loss, achieving tracking performance nearly identical to full-episode results.
@@ VI-A Data Collection
We collect one GelSight video per object for qualitative analysis on a set of 15 real-world objects, including $3$ tool handles, $7$ common food items, $4$ rocks and fossils, and $1$ textured object. The smallest object is Seed ($8\,mm\times 8\,mm\times 8\,mm$), and the largest is Avocado ($85\,mm\times 61\,mm\times 58\,mm$). All objects have at least some surface texture, with the smoothest being the handle of a wooden spoon. Video durations range from $1$ to $30$ minutes. Fig. 15 shows each object with its size and video duration. For quantitative analysis, we use 5 object models with known ground-truth geometries. Each model is 3D-printed in two sizes using a Form 3+ printer, with a reported average dimensional deviation of up to 0.1 mm, resulting in a total of 10 printed objects. The smallest and largest objects are a Seed and Lime with a diameter of about 9 mm and 30 mm, respectively. All data are collected in an “in-the-wild” setup, where we manually hold both the object and the GelSight Mini sensor (Fig. 12) during scanning. Contact breaks and re-initializations are common during data collection, sometimes exceeding 100 occurrences per object. Scanning paths are guided by real-time reconstruction visualization, which the operator uses to adjust sensor movement during scanning. For very long scans, we segment data collection into consecutive 10-minute scans and fuse the resulting tactile videos into a single sequence for reconstruction.
@@ VI-B System Characteristics and Performance
We run GelSLAM in offline mode on all object videos. We focus on offline reconstruction here, as offline shape reconstruction is sufficient for most reconstruction applications, and discuss online reconstruction in Section VII-B. Since the objects provide sufficient texture, we apply fixed and less conservative thresholds for NormalFlow failure detection: the estimate is rejected if $\text{CCS}<0.7$ or $\text{SCR}<0.3$ across all objects. Table V summarizes reconstruction statistics for four objects with varying video lengths, and Fig. 13 shows their corresponding pose graphs.
@@
Object Name Avocado Large Rock Peanut Seed
@@
GelSLAM Offline Runtime (min) 40.92 29.82 4.78 0.6
@@ VI-C Qualitative Results
Fig. 15 shows tactile-based object-level 3D reconstructions for the $15$ real-world objects. All loops detected by GelSLAM are correct, indicating its robustness. The results demonstrate that GelSLAM correctly relates GelSight images spatially and globally reconstructs objects across a wide range of textures and sizes, even on low-texture objects such as the wooden spoon handle. The reconstructed geometry also exhibits remarkable detail.